The radius of a coil of wire with N turns is r=0.03m. A counter clockwise current of 1Amp flows through the coil. At a point is carrying a current=4 Amps and is located 0.005m from the center of the coil.(its inside the coil) Determine N, the number of turns in the Coil.
Answer #1

we know on the axis of circular coil, B = N*mue*I*R^2/(2*(x^2+R^2)^(3/2))

so, N = B*(2*(x^2+R^2)^(3/2))/(mue*I*R^2)

= 8.136*(2*(0.005^2 + 0.03^2)^1.5)/(4*pi*10^-7*4*0.03^2)

= 101191 turns